# A Watermark for Low-entropy and Unbiased Generation in Large Language Models

### v2

@@ Appendix E STA-M Details
$p_{a}^{w,k}=\begin{cases}\frac{1}{2}+\frac{1}{6}\times\frac{1}{2}+(\frac{1}{6})^{2}\times\frac{1}{2}+\cdots+(\frac{1}{6})^{M}\times\frac{1}{2}&\{a,b\in G;c,d\in R\}\\ \frac{1}{2}+\frac{5}{12}\times\frac{1}{2}+(\frac{5}{12})^{2}\times\frac{1}{2}+\cdots+(\frac{5}{12})^{M}\times\frac{1}{2}&\{a,c\in G;b,d\in R\}\\ \frac{1}{2}+\frac{5}{12}\times\frac{1}{2}+(\frac{5}{12})^{2}\times\frac{1}{2}+\cdots+(\frac{5}{12})^{M}\times\frac{1}{2}&\{a,d\in G;b,c\in R\}\\ (\frac{7}{12})^{M}\times\frac{1}{2}&\{b,c\in G;a,d\in R\}\\ (\frac{7}{12})^{M}\times\frac{1}{2}&\{b,d\in G;a,c\in R\}\\ (\frac{5}{6})^{M}\times\frac{1}{2}&\{c,d\in G;a,b\in R\}\\ \end{cases}.$
@@
$\mathbb{E}_{k\sim P_{K}(k)}\left[p_{a}^{w,k}\right]$ $=\frac{1}{12}\left[\frac{6}{5}\left(1-(\frac{1}{6})^{M+1}\right)+2\times\frac{12}{7}\left(1-(\frac{5}{12})^{M+1}\right)+2\times(\frac{7}{12})^{M}+(\frac{5}{6})^{M}\right]$
$=\frac{27}{70}-\frac{1}{10}(\frac{1}{6})^{M+1}-\frac{2}{7}(\frac{5}{12})^{M+1}+\frac{1}{6}(\frac{7}{12})^{M}+\frac{1}{12}(\frac{5}{6})^{M},$
@@ F.1 Experimental Setup
(a) STA-4 on C4
(b) STA-8 on C4
(c) STA-16 on C4
(d) STA-4 on HumanEval
(e) STA-8 on HumanEval
(f) STA-16 on HumanEval
Figure 1: Performance of STA-M w.r.t. $\tau$
Table 6: Examples of STA-generated Texts for C4
Prompt Human-written STA-1 generated STA-16 generated
[]delimited-[][$\dots$][ … ] Thomas will be responsible for overseeing Micron’s solid state storage business that ranges from hard disk drive replacements with solid state drives (SSDs) to enterprise-class storage solutions. He brings more than 30 years of experience to Micron and most recently served as the vice president of Enterprise Storage for Micron’s common stock is traded on the NASDAQ under the MU symbol. To learn more about Micron Technology, Inc., visit www.micron.com. Micron and the Micron orbit logo are trademarks of Micron Technology, Inc. All other trademarks are the property of their respective owners. []delimited-[][$\dots$][ … ] (PPL:3.25) the Americas region for Seagate Technology. He is a senior executive level leader with a proven track record in defining strategy that drives revenue, profit and new technology execution. ”Micron is thrilled to have Darren as part of our team,” said Mary Jane Raymond, . []delimited-[][$\dots$][ … ] (PPL:4.45) Fusion I/O, LLC. Before that, Thomas was at Western Digital Corporation where he was a progressive executive, holding various management roles since 2008, most recently as its executive vice president of storage technology. []delimited-[][$\dots$][ … ]
Table 7: Examples of STA-generated Texts for HumanEval
Prompt Canonical Solution STA-1 generated STA-16 generated
from typing import List def filter_by_prefix (strings: List[str], prefix: str) -$>$ List[str]: ””” Filter an input list of strings only for ones that start with a given prefix. $>$much-greater-thanabsent>>>> > > filter_by_prefix([], ’a’) [] $>$much-greater-thanabsent>>>> > > filter_by_prefix([’abc’, ’bcd’, ’cde’, ’array’], ’a’) [’abc’, ’array’] ””” return [x for x in strings if x.startswith(prefix)] (Pass, PPL:1.92) from typing import List def filter_by_prefix (strings: List[str], prefix: str) -$>$ List[str]: ””” Filter an input list of strings only for ones that start with a given prefix. $>$much-greater-thanabsent>>>> > > filter_by_prefix([], ’a’) [] $>$much-greater-thanabsent>>>> > > filter_by_prefix([’abc’, ’bcd’, ’cde’, ’array’], ’a’) [’array’, ’abc’] ””” return [word for word in strings if word.startswith(prefix)] (Pass, PPL:2.18) from typing import List def filter_by_prefix (strings: List[str], prefix: str) -$>$ List[str]: ”””* Filter an input list of strings only for ones that start with a given prefix. Input: ‘strings = [”abc”, ”bcd”, ”cde”, ”array”]‘, ‘prefix = ’a’‘ Output: ‘[”abc”, ”array”]‘ ””” return [word for word in strings if word.startswith(prefix)]
from typing import List def intersperse (numbers: List[int], delimeter: int) -$>$ List[int]: ””” Insert a number ’delimeter’ between every two consecutive elements of input list ‘numbers’ $>$much-greater-thanabsent>>>> > > intersperse([], 4) [] $>$much-greater-thanabsent>>>> > > intersperse([1, 2, 3], 4) [1, 4, 2, 4, 3] ””” if not numbers: return [] result = [] for n in numbers[:-1]: result.append(n) result.append(delimeter) result.append(numbers[-1]) return result (Not pass, PPL:2.45) def intersperse (numbers: List[int], delimeter: int) -$>$ List[int]: ””” Insert a number ’delimeter’ between every two consecutive elements of input list ‘numbers’ $>$much-greater-thanabsent>>>> > > intersperse([], 4) [] $>$much-greater-thanabsent>>>> > > intersperse([1, 2, 3], 4) [1, 4, 2, 4, 3] ””” answer=[] for i in range(0,len(numbers),1): answer = answer + [ numbers[i] ] answer = answer + [ delimeter ] return answer (Not pass, PPL:3.38) Maybe you meant () -$>$ List[int]
def is_prime(n): ”””Return true if a given number is prime, and false otherwise. ”’ if n ¡ 2: return False for k in range(2, n - 1): if n % k == 0: return False return True (Pass, PPL:1.94) “‘python def is_prime(n): if n == 1: return False for i in range(2, n): if n % i == 0: return False return True (Pass, PPL:1.78) def is_prime(n): if n == 1: return False for i in range(2, n): if n % i == 0: return False return True if __name__ == ’__main__’: import doctest doctest.testmod()
@@ F.2 Robustness Check on Entropy Threshold Parameter
In this section, we conducted a robustness check on the parameter $\tau$ in STA-M. In particular, we set the low entropy threshold $\tau$ from 0 to 2.1 with an interval of 0.15. At each generation step, we sampled at most 4, 8, and 16 times (i.e., STA-4, STA-8, and STA-16) when the entropy was above the threshold $\tau$. Figure 1 shows text quality and watermark strength of STA-M with different $\tau$s. As depicted, different $\tau$s do not affect the watermark strength significantly for C4 because C4 is a high-entropy dataset. Also, we observe a decrease in PPL when we increase $\tau$ in Figure 1(a), 1(b), and 1(c). The reason is that by setting up a higher entropy threshold, fewer generation steps will apply the STA-M strategy, making the watermarking method more similar to STA-1. According to Figure 1(d), 1(e), and 1(f), we observe a general increase of watermark strength if we have a larger $\tau$ because we will have more green list tokens if we sample M times instead of once. However, higher watermark strength leads to a lower pass@$1$ score, which is related to the text quality (Kirchenbauer et al., 2023a). We chose the Pareto optimal of each dataset as our final parameter for each dataset. Specifically, we selected $\tau=1.35$ for C4 and $\tau=1.95$ for HumanEval.

### v3

@@ Appendix E STA-M Details
$p_{a}^{w,k}=\begin{cases}\frac{1}{2}+\frac{1}{6}\times\frac{1}{2}+\cdots+(\frac{1}{6})^{M}\times\frac{1}{2}\\ (\{a,b\in G;c,d\in R\})\\ \frac{1}{2}+\frac{5}{12}\times\frac{1}{2}+\cdots+(\frac{5}{12})^{M}\times\frac{1}{2}\\ (\{a,c\in G;b,d\in R\})\\ \frac{1}{2}+\frac{5}{12}\times\frac{1}{2}+\cdots+(\frac{5}{12})^{M}\times\frac{1}{2}\\ (\{a,d\in G;b,c\in R\})\\ (\frac{7}{12})^{M}\times\frac{1}{2}\\ (\{b,c\in G;a,d\in R\})\\ (\frac{7}{12})^{M}\times\frac{1}{2}\\ (\{b,d\in G;a,c\in R\})\\ (\frac{5}{6})^{M}\times\frac{1}{2}\\ (\{c,d\in G;a,b\in R\})\\ \end{cases}.$
@@
$\mathbb{E}_{k\sim P_{K}(k)}\left[p_{a}^{w,k}\right]$
$=\frac{1}{12}\bigg{[}\frac{6}{5}\left(1-(\frac{1}{6})^{M+1}\right)$
$\quad+2\times\frac{12}{7}\left(1-(\frac{5}{12})^{M+1}\right)$
$\quad+2\times(\frac{7}{12})^{M}+(\frac{5}{6})^{M}\bigg{]}$
$=\frac{27}{70}-\frac{1}{10}(\frac{1}{6})^{M+1}-\frac{2}{7}(\frac{5}{12})^{M+1}$
$\quad+\frac{1}{6}(\frac{7}{12})^{M}+\frac{1}{12}(\frac{5}{6})^{M},$
@@
(a) STA-4 on C4
(b) STA-8 on C4
(c) STA-16 on C4
(d) STA-4 on HumanEval
(e) STA-8 on HumanEval
(f) STA-16 on HumanEval
Figure 4: Performance of STA-M w.r.t. $\tau$
Table 4: Examples of STA-generated Texts for C4
Prompt Human-written STA-1 generated STA-16 generated
[]delimited-[][$\dots$][ … ] Thomas will be responsible for overseeing Micron’s solid state storage business that ranges from hard disk drive replacements with solid state drives (SSDs) to enterprise-class storage solutions. He brings more than 30 years of experience to Micron and most recently served as the vice president of Enterprise Storage for Micron’s common stock is traded on the NASDAQ under the MU symbol. To learn more about Micron Technology, Inc., visit www.micron.com. Micron and the Micron orbit logo are trademarks of Micron Technology, Inc. All other trademarks are the property of their respective owners. []delimited-[][$\dots$][ … ] (PPL:3.25) the Americas region for Seagate Technology. He is a senior executive level leader with a proven track record in defining strategy that drives revenue, profit and new technology execution. "Micron is thrilled to have Darren as part of our team," said Mary Jane Raymond, . []delimited-[][$\dots$][ … ] (PPL:4.45) Fusion I/O, LLC. Before that, Thomas was at Western Digital Corporation where he was a progressive executive, holding various management roles since 2008, most recently as its executive vice president of storage technology. []delimited-[][$\dots$][ … ]
Table 5: Examples of STA-generated Texts for HumanEval
Prompt Canonical Solution STA-1 generated STA-16 generated
from typing import List def filter_by_prefix (strings: List[str], prefix: str) -$>$ List[str]: """ Filter an input list of strings only for ones that start with a given prefix. $>$much-greater-thanabsent>>>> > > filter_by_prefix([], ’a’) [] $>$much-greater-thanabsent>>>> > > filter_by_prefix([’abc’, ’bcd’, ’cde’, ’array’], ’a’) [’abc’, ’array’] """ return [x for x in strings if x.startswith(prefix)] (Pass, PPL:1.92) from typing import List def filter_by_prefix (strings: List[str], prefix: str) -$>$ List[str]: """ Filter an input list of strings only for ones that start with a given prefix. $>$much-greater-thanabsent>>>> > > filter_by_prefix([], ’a’) [] $>$much-greater-thanabsent>>>> > > filter_by_prefix([’abc’, ’bcd’, ’cde’, ’array’], ’a’) [’array’, ’abc’] """ return [word for word in strings if word.startswith(prefix)] (Pass, PPL:2.18) from typing import List def filter_by_prefix (strings: List[str], prefix: str) -$>$ List[str]: """* Filter an input list of strings only for ones that start with a given prefix. Input: ‘strings = ["abc", "bcd", "cde", "array"]‘, ‘prefix = ’a’‘ Output: ‘["abc", "array"]‘ """ return [word for word in strings if word.startswith(prefix)]
from typing import List def intersperse (numbers: List[int], delimeter: int) -$>$ List[int]: """ Insert a number ’delimeter’ between every two consecutive elements of input list ‘numbers’ $>$much-greater-thanabsent>>>> > > intersperse([], 4) [] $>$much-greater-thanabsent>>>> > > intersperse([1, 2, 3], 4) [1, 4, 2, 4, 3] """ if not numbers: return [] result = [] for n in numbers[:-1]: result.append(n) result.append(delimeter) result.append(numbers[-1]) return result (Not pass, PPL:2.45) def intersperse (numbers: List[int], delimeter: int) -$>$ List[int]: """ Insert a number ’delimeter’ between every two consecutive elements of input list ‘numbers’ $>$much-greater-thanabsent>>>> > > intersperse([], 4) [] $>$much-greater-thanabsent>>>> > > intersperse([1, 2, 3], 4) [1, 4, 2, 4, 3] """ answer=[] for i in range(0,len(numbers),1): answer = answer + [ numbers[i] ] answer = answer + [ delimeter ] return answer (Not pass, PPL:3.38) Maybe you meant () -$>$ List[int]
def is_prime(n): """Return true if a given number is prime, and false otherwise. ”’ if n < 2: return False for k in range(2, n - 1): if n % k == 0: return False return True (Pass, PPL:1.94) “‘python def is_prime(n): if n == 1: return False for i in range(2, n): if n % i == 0: return False return True (Pass, PPL:1.78) def is_prime(n): if n == 1: return False for i in range(2, n): if n % i == 0: return False return True if __name__ == ’__main__’: import doctest doctest.testmod()
Table 6: Attacking Watermarks for the C4 Dataset.
        Attack Setting No Attack Copy-Paste GPT-3.5 DIPPER-1 DIPPER-2
Method $\uparrow$ F1 $\uparrow$ AUC $\uparrow$ F1 $\uparrow$ AUC $\uparrow$ F1 $\uparrow$ AUC $\uparrow$ F1 $\uparrow$ AUC $\uparrow$ F1 $\uparrow$ AUC
RDW 0.98 0.98 0.77 0.79 0.43 0.62 0.34 0.53 0.45 0.63
Dipmark($\alpha=0.3$) 0.93 0.94 0.61 0.70 0.29 0.57 0.24 0.55 0.26 0.55
Dipmark($\alpha=0.4$) 0.96 0.96 0.75 0.79 0.38 0.61 0.31 0.58 0.34 0.59
$\gamma$-reweight 0.96 0.96 0.74 0.78 0.41 0.61 0.32 0.57 0.36 0.60
STA-1 0.96 0.96 0.78 0.81 0.47 0.63 0.39 0.60 0.46 0.63
KGW($\delta=1$) 0.96 0.96 0.68 0.75 0.27 0.57 0.13 0.53 0.15 0.54
KGW($\delta=1.5$) 0.99 0 98 0.90 0.90 0.41 0.62 0.22 0.56 0.27 0.57
KGW($\delta=2$) 0.99 0.99 0.95 0.95 0.54 0.68 0.30 0.58 0.40 0.62
STA-4($\tau$=1.35) 0.97 0.97 0.95 0.95 0.72 0.78 0.65 0.73 0.69 0.75
STA-8($\tau$=1.35) 0.98 0.98 0.95 0.95 0.78 0.81 0.71 0.77 0.76 0.79
STA-16($\tau$=1.35) 0.97 0.97 0.95 0.95 0.76 0.80 0.68 0.74 0.78 0.81
@@ F.2 Robustness Check on Entropy Threshold Parameter
In this section, we conducted a robustness check on the parameter $\tau$ in STA-M. In particular, we set the low entropy threshold $\tau$ from 0 to 2.1 with an interval of 0.15. At each generation step, we sampled at most 4, 8, and 16 times (i.e., STA-4, STA-8, and STA-16) when the entropy was above the threshold $\tau$. Figure 4 shows text quality and watermark strength of STA-M with different $\tau$s. As depicted, different $\tau$s do not affect the watermark strength significantly for C4 because C4 is a high-entropy dataset. Also, we observe a decrease in PPL when we increase $\tau$ in Figure 4(a), 4(b), and 4(c). The reason is that by setting up a higher entropy threshold, fewer generation steps will apply the STA-M strategy, making the watermarking method more similar to STA-1. According to Figure 4(d), 4(e), and 4(f), we observe a general increase of watermark strength if we have a larger $\tau$ because we will have more green list tokens if we sample M times instead of once. However, higher watermark strength leads to a lower pass@$1$ score, which is related to the text quality (Kirchenbauer et al., 2023a). We chose the Pareto optimal of each dataset as our final parameter for each dataset. Specifically, we selected $\tau=1.35$ for C4 and $\tau=1.95$ for HumanEval.
@@ F.4 Attacking Watermarks
For the copy-paste attack, since STA-1 and STA-M are based on the green-red list partition and changing a token can only affect the detection score of itself and the next token, it is naturally robust to simple text insertion and removal (Kirchenbauer et al., 2023a). Meanwhile, LLM-based attacks, such as GPT-3.5 and DIPPER, are designed to replace tokens in given texts by sampling from the LLM. STA-M effectively increases the proportion of green-list tokens by raising their probability in high-entropy scenarios without compromising too much text quality, making it difficult for LLM-based attacks to replace a substantial number of tokens in STA-M-generated text and remove the watermark.